Circular No. 7757

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(90) ANTIOPE T. Michałowski, T. Kwiatkowski, and A. Kryszczyńska, Poznań As-

tronomical Observatory; F. Colas, L’Institut de Mécanique Céleste et de Calcul des Éphémérides, Paris; and J. Michałowski, Poznań University of Technology, write: “CCD observations of the binary minor planet (90) Antiope (cf. IAUC 7503), carried out at Borowiec Observatory and at Pic du Midi on seven nights between Oct. 19 and Nov. 15, show a two-component lightcurve with each showing the same period of 16.496 hr. The first component (with amplitude 0.10 mag) is associated with the rotation of two non-spherical bodies of the system; the second one, showing two sharp minima (with amplitude 0.12 mag), is due to mutual occultation/eclipse events in the binary system. The lightcurve suggests that the rotational periods of both bodies are equal to the orbital period, which is characteristic for synchronous rotation. The amplitude of the existing occultation/eclipse is much larger than predicted earlier by Michałowski et al. (2001, A.Ap. 378, L14), which suggests that the bodies are larger than previously thought and/or that the orbital plane of the system precesses. New observations will help to improve a model of the binary minor planet.”

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD R. N. Manchester, Australia Telescope National Facility; and B. M.

Gaensler, Harvard-Smithsonian Center for Astrophysics, report on behalf of a larger collaboration: “We have observed SN 1987A with the newly installed 12-mm receivers on the Australia Telescope Compact Array. In a 12-hr observation beginning Oct. 26.59 UT, we detected the supernova remnant at a high significance level. Its position is consistent with previous radio observations (Gaensler et al. 1997, Ap.J. 479, 845; Manchester et al. 2001, Pub. Astron. Soc. Australia, in press). The observed flux densities were 20 ± 1 and 18 ± 1 mJy at 17.0 and 18.9 GHz, respectively. These measurements are marginally above extrapolations of the power-law spectrum previously seen between 0.8 GHz and 9 GHz. The source is slightly extended at our available spatial resolution (2′′.8 × 1′′.8). Modeling the data as a thin spherical shell yields a diameter of 1′′.6 ± 0′′.1, consistent with the slow expansion speed inferred from lower frequency measurements. Further 12-mm observations at higher resolution will be carried out later this year.”

2001 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green